# Soot Planets instead of Water Worlds

### v1

@@ Appendix D M-R Relations
We calculated the mass-radius (M-R) relations of model planets through a finite-element iterative algorithm, assuming the Rose-Vinet form of equation-of-state for all phases at 300 K, except hydrogen (P. Vinet et al., 1989),
@@
The thickness of a hydrogen envelope surrounding a rocky planet is calculated using a fitted polynomial expression for hydrogen across gaseous and solid states (M. Tkacz & A. Litwiniuk, 2002), instead of a Vinet EoS for crystalline hydrogen (e.g., P. Loubeyre et al., 1996) that is only applicable at pressures above $\sim$10 GPa.
@@
aAll values are at ambient conditions (100 kPa and 298 K), except that the density of CH4 is at 100 kPa and 30 K (T. P. Mangan et al., 2017). Uncertainties are on the last digits, if not listed.  bAssuming the densities of Febcc (T. Komabayashi & Y. Fei, 2010), forsterite (G. J. Finkelstein et al., 2014), graphite (Y. Wang et al., 2012), and liquid H2O (V. B. Prakapenka et al., 2021), the stable phases at ambient conditions.  cFitted polynomial $V=AP^{-\frac{1}{3}}+BP^{-\frac{2}{3}}+CP^{-\frac{4}{3}}+(D+ET)P^{-1}$, where $P$ is pressure in GPa, and $T$ is temperature in K (M. Tkacz & A. Litwiniuk, 2002).  Data sources: [1] hcp = hexagonal close packed R. F. Smith et al. (2018); [2] bdg = bridgmanite Y. Fei et al. (2021); [3] dia = diamond D. K. Bradley et al. (2009); [4] Applies to ice VII, VII’, and X V. B. Prakapenka et al. (2021); [5] fcc = face centered cubic T. Komabayashi & Y. Fei (2010); [6] ol = olivine (forsterite) G. J. Finkelstein et al. (2014); [7] gra = graphite Y. Wang et al. (2012); [8] L. Sun et al. (2009). [9] pm = paramagnetic B. Chen et al. (2012); [10] ppv = post-perovskite S. R. Shieh et al. (2006); [11] SI = super-ionic fluid V. B. Prakapenka et al. (2021); [12] P. Loubeyre et al. (1996).
@@
Figure A4: (A) The compression curves of iron at 300 K span a range that encloses the isentropic path of iron (R. F. Smith et al., 2018). The vertical bar marks the density jump across the bcc–hcp transition at 14 GPa and 300 K. (B) Comparison of isothermal compression curves of MgSiO3 at 300 K and 7800 K (Y. Fei et al., 2021). Vertical bars mark the density jumps across the perovskite (pv) to post-perovskite (ppv) transition, and the ppv to ppv1 transition. (C) Comparison of 300 K compression curves (D. C. Swift et al., 2022; R. F. Smith et al., 2018) and high-temperature compression curve (R. F. Smith et al., 2018) of diamond. In one study (R. F. Smith et al., 2018), the 300 K range (gray) overlaps with the high-temperature (high-$T$) curve. (D) Compression curves of H2O at 300 K (this study), compared with isothermal curves at 1000 K and 6000 K (M. French et al., 2009), and with density–pressure profiles from the interiors of Neptune and hot Neptunes (M. D. Knudson et al., 2012). (E) Comparison of hydrogen equations of state (EoS). The empirical TEoS from metallurgical literature (M. Tkacz & A. Litwiniuk, 2002) can be used to calculate the mass–radius relation of planets with H2 atmospheres. Although not designed for extreme pressures, the polynomial fit reproduces tabulated H-REOS.3 data (A. Becker et al., 2014) (based on ab initio calculations for astrophysical use). Discrepancies between the two EoS sets are negligible above 20 GPa, but grow at low volumes (high pressures >>100 GPa) due to the inverse relation between density and volume.
@@
The interiors of planets are expected to be hot due to adiabatic compression and partial conversion of the energy of accretion and differentiation into heat. Phase transitions and chemical reactions are common under the pressure and temperature conditions of planetary interiors (Fig. A4). For example, methane (CH4) undergoes a series of phase transitions at high pressures and 300 K (L. Sun et al., 2009). Theory predicts that methane dissociates into ethane (C2H6) at 95 GPa, then butane (C4H10) at 158 GPa, and further, diamond (C) and hydrogen (H2) above 287 GPa at 0 K (G. Gao et al., 2010).
@@
The internal thermal profiles of exoplanets are highly uncertain, as they depend on the complex processes of accretion and early differentiation, the retention factors of formation energies, as well as the composition, long-term evolution, and age of planets ((N. I. White & J. Li, 2025) and references therein).
@@
There are limited constraints on material properties under relevant conditions (Table A3). Existing data suggest that the influences of phase changes and thermal expansion on densities are often smaller than the uncertainties in the 300-K compression curves at high pressures(Fig. A4). Fortunately, thermal expansion decreases rapidly with pressure, and therefore the expected effects are limited both at low pressures where temperatures are relatively low, and at high pressures where the thermal expansion coefficient is small. Phase transitions could lead to densification or density reduction. While pressure-induced transitions into denser structures partially offset that of thermal expansion on densities, heating would induce transitions into phases of lower densities. In addition, material structure at very high pressure is mainly supported by the electron degeneracy pressure, which has similar functional dependence on compression for different materials, and therefore the high-pressure density chiefly depends on the mean atomic number and secondarily on the crystal structure or temperature (L. Zeng et al., 2019).
@@
Previous studies considered phase and temperature changes for low-mass planet models (e.g., D. Valencia et al., 2007; C. Sotin et al., 2007), Although more realistic, they involve complicated interior boundary conditions and highly uncertain EoS, and yielded M-R relations that agree with our simplified models (Fig. A5).
@@
Figure A5: Comparison of mass–radius relations for iron planets, Earth-like rocky planets, and water-worlds composed of 50% rock and 50% water. Curves from models that include phase transitions and thermal effects are shown in red (dotted: (L. Zeng et al., 2019); short-dashed: (D. Valencia et al., 2007); long-dashed: (C. Sotin et al., 2007); solid: (R. F. Smith et al., 2018)). Models computed at 300 K are shown in black (solid: this study; dashed: (R. F. Smith et al., 2018)).

### v2

@@ Appendix D M-R Relations
We calculated the mass-radius (M-R) relations of model planets through a finite-element iterative algorithm, assuming the Rose-Vinet form of equation-of-state for all phases at 300 K, except hydrogen (Vinet1989),
@@
A multi-layer planet consists of four distinct layers – an iron core, a silicate mantle, a soot layer, and an outer water layer. A single-layer planet is modeled as a uniform mixture of iron, silicate, soot, and ice, using the Voigt-Reuss-Hill average:
$\rho_{v}$ $=\sum_{i}m_{i}\rho_{i}$
$\rho_{r}$ $=\frac{1}{\sum_{i}\frac{m_{i}}{\rho_{i}}}$
$K_{v}$ $=\sum_{i}m_{i}K_{i}$
$K_{r}$ $=\frac{1}{\sum_{i}\frac{m_{i}}{K_{i}}}$
$Kp_{v}$ $=\sum_{i}m_{i}Kp_{i}$
$Kp_{r}$ $=\frac{1}{\sum_{i}\frac{m_{i}}{Kp_{i}}}$
$\rho$ $=\tfrac{1}{2}(\rho_{v}+\rho_{r})$
$K$ $=\tfrac{1}{2}(K_{v}+K_{r})$
$Kp$ $=\tfrac{1}{2}(Kp_{v}+Kp_{r})$
where $i$ is the index over components, $m_{i}$ is the mass fraction of component $i$ (with $\sum_{i}m_{i}=1$), $\rho_{i}$ is the density of component $i$, $K_{i}$ is the bulk modulus of component $i$, $Kp_{i}$ is the pressure derivative of the bulk modulus of component $i$, $\sum_{i}$ denotes summation over all components, $\rho_{v}$ is the Voigt average density (arithmetic mean), $\rho_{r}$ is the Reuss average density (harmonic mean), $K_{v}$ is the Voigt average bulk modulus, $K_{r}$ is the Reuss average bulk modulus, $Kp_{v}$ is the Voigt average of $Kp$, $Kp_{r}$ is the Reuss average of $Kp$, $\rho$ is the Hill averaged density, $K$ is the Hill averaged bulk modulus, $Kp$ is the Hill averaged pressure derivative, and $\tfrac{1}{2}$ indicates taking the arithmetic mean of the Voigt and Reuss bounds. We note this is an approximation as a homogeneous phase formed by chemically miscible components is not equivalent to a physical mixture of distinct phases.
The thickness of a hydrogen envelope surrounding a rocky planet is calculated using a fitted polynomial expression for hydrogen across gaseous and solid states (Tkacz2002), instead of a Vinet EoS for crystalline hydrogen (e.g., Loubeyre1996) that is only applicable at pressures above $\sim$10 GPa.
@@
aAll values are at ambient conditions (100 kPa and 298 K), except that the density of CH4 is at 100 kPa and 30 K (Mangan2017). Uncertainties are on the last digits, if not listed.  bAssuming the densities of Febcc (komabayashi2010), forsterite (Finkelstein14Forsterite), graphite (Wang12Graphite), and liquid H2O (Prakapenka21), the stable phases at ambient conditions.  cFitted polynomial $V=AP^{-\frac{1}{3}}+BP^{-\frac{2}{3}}+CP^{-\frac{4}{3}}+(D+ET)P^{-1}$, where $P$ is pressure in GPa, and $T$ is temperature in K (Tkacz2002).  Data sources: [1] hcp = hexagonal close packed Smith2018_fe_eos; [2] bdg = bridgmanite Fei2021; [3] dia = diamond Bradley2009; [4] Applies to ice VII, VII’, and X Prakapenka21; [5] fcc = face centered cubic komabayashi2010; [6] ol = olivine (forsterite) Finkelstein14Forsterite; [7] gra = graphite Wang12Graphite; [8] Sun2009. [9] pm = paramagnetic Chen12; [10] ppv = post-perovskite Shieh06ppv; [11] SI = super-ionic fluid Prakapenka21; [12] Loubeyre1996.
@@
Figure A4: (A) The compression curves of iron at 300 K span a range that encloses the isentropic path of iron (Smith2018_fe_eos). The vertical bar marks the density jump across the bcc–hcp transition at 14 GPa and 300 K. (B) Comparison of isothermal compression curves of MgSiO3 at 300 K and 7800 K (Fei2021). Vertical bars mark the density jumps across the perovskite (pv) to post-perovskite (ppv) transition, and the ppv to ppv1 transition. (C) Comparison of 300 K compression curves (2022arXiv220316065S; Smith2018_fe_eos) and high-temperature compression curve (Smith2018_fe_eos) of diamond. In one study (Smith2018_fe_eos), the 300 K range (gray) overlaps with the high-temperature (high-$T$) curve. (D) Compression curves of H2O at 300 K (this study), compared with isothermal curves at 1000 K and 6000 K (2009PhRvB..79e4107F), and with density–pressure profiles from the interiors of Neptune and hot Neptunes (2012PhRvL.108i1102K). (E) Comparison of hydrogen equations of state (EoS). The empirical TEoS from metallurgical literature (Tkacz2002) can be used to calculate the mass–radius relation of planets with H2 atmospheres. Although not designed for extreme pressures, the polynomial fit reproduces tabulated H-REOS.3 data (2014ApJS..215...21B) (based on ab initio calculations for astrophysical use). Discrepancies between the two EoS sets are negligible above 20 GPa, but grow at low volumes (high pressures >>100 GPa) due to the inverse relation between density and volume.
@@
The interiors of planets are expected to be hot due to adiabatic compression and partial conversion of the energy of accretion and differentiation into heat. Phase transitions and chemical reactions are common under the pressure and temperature conditions of planetary interiors (Fig. A4). For example, methane (CH4) undergoes a series of phase transitions at high pressures and 300 K (Sun2009). Theory predicts that methane dissociates into ethane (C2H6) at 95 GPa, then butane (C4H10) at 158 GPa, and further, diamond (C) and hydrogen (H2) above 287 GPa at 0 K (Gao2010methane).
@@
The internal thermal profiles of exoplanets are highly uncertain, as they depend on the complex processes of accretion and early differentiation, the retention factors of formation energies, as well as the composition, long-term evolution, and age of planets ((White25) and references therein).
@@
There are limited constraints on material properties under relevant conditions (Table A3). Existing data suggest that the influences of phase changes and thermal expansion on densities are often smaller than the uncertainties in the 300-K compression curves at high pressures(Fig. A4). Fortunately, thermal expansion decreases rapidly with pressure, and therefore the expected effects are limited both at low pressures where temperatures are relatively low, and at high pressures where the thermal expansion coefficient is small. Phase transitions could lead to densification or density reduction. While pressure-induced transitions into denser structures partially offset that of thermal expansion on densities, heating would induce transitions into phases of lower densities. In addition, material structure at very high pressure is mainly supported by the electron degeneracy pressure, which has similar functional dependence on compression for different materials, and therefore the high-pressure density chiefly depends on the mean atomic number and secondarily on the crystal structure or temperature (Zeng2019).
@@
Previous studies considered phase and temperature changes for low-mass planet models (e.g., Valencia2007ApJ...665.1413V; Sotin2007Icar..191..337S), Although more realistic, they involve complicated interior boundary conditions and highly uncertain EoS, and yielded M-R relations that agree with our simplified models (Fig. A5).
@@
Figure A5: Comparison of mass–radius relations for iron planets, Earth-like rocky planets, and water-worlds composed of 50% rock and 50% water. Curves from models that include phase transitions and thermal effects are shown in red (dotted: (Zeng2019); short-dashed: (Valencia2007ApJ...665.1413V); long-dashed: (Sotin2007Icar..191..337S); solid: (Smith2018_fe_eos)). Models computed at 300 K are shown in black (solid: this study; dashed: (Smith2018_fe_eos)).